# Semantic-Aware Spectrum Sharing in Internet of Vehicles Based on Deep Reinforcement Learning

### v3

@@ II-C Optimization Problem
${P_{0}}:\mathop{\max}\limits_{{u_{q}}}\sum\limits_{w=1}^{W}{HSSE_{q,w}^{VI}}+\sum\limits_{w=1}^{W}{\sum\limits_{q=1}^{Q}{{b_{q}}\left[w\right]}}p_{q}^{VV}\left[w\right]$ (15a)
@@
Constraints (15b), (15c) and (15d) represent the channel allocation. (15c) states that each V2V link can only share one of the V2I links, (15d) specifies that each V2I link can only be allocated to one V2V link, (15e) constrains the average number of semantic symbols for each word within a certain range, and (15f) constrains the minimum value of semantic similarity.
@@ III-A2 Action
The key actions include channel selection for V2V links, transmission power control, and the selection of the length of semantic symbols ${u_{q}}$ by the V2V transmitter vehicle in the shared V2I link. The agent’s actions can be represented as a vector containing multiple elements. Specifically, it can be represented as shown in (17):
${a_{q}}\left(t\right)=\left({{B_{q}}\left(t\right),{P_{q}}\left(t\right),{L_{q}}\left(t\right)}\right),$ (17)
where ${B_{q}}\left(t\right)$ represents the channel selection, ${P_{q}}\left(t\right)$ represents the transmission power chosen by the V2V link transmitter vehicles, and ${L_{q}}\left(t\right)$ represents the length of semantic symbols chosen by the V2V transmitter vehicle in the shared V2I link. These actions’ values are usually confined within the range $\left[{-1,0}\right]$, which has been selected for efficient learning and convergence. Subsequently, these values undergo a linear mapping process to obtain actual parameters or lengths. Defining this action space enables flexible decision-making in semantically-aware spectrum sharing within high-speed mobile IoV environments, aiming to maximize the HSSE and SRS over V2V links.
@@ III-B SSS Algorithm
The SSS algorithm regarding the training phase is illustrated in Algorithm 1 and is described in detail as follows.
@@ III-B1 Training stage
Firstly, the parameters $\rho$, ${\theta_{1}}$, and ${\theta_{2}}$, along with the weight parameters $\varepsilon$ and the discounted factor $\gamma$, are randomly initialized. Set ${\bar{\theta}_{1}}$ and ${\bar{\theta}_{2}}$ equal to ${\theta_{1}}$ and ${\theta_{2}}$, respectively. Initialize the mobile vehicle network in the urban environment, specifying the number and positions of vehicles, as well as the values for fast and slow channel fading.
@@
where $sh{a_{new}}$ and $sh{a_{old}}$ is the shadow fading at the new and old position, respectively. $X\sim N(0,{3^{2}})$ , ${\Delta d}$ represents the distance between paths. ${{d_{dec}}}$ is the decorrelation distance used to describe the correlation of shadow fading.
@@
Next, initialize the size of the semantic data packets to be transmitted, the specified transmission time, and the link.
The algorithm iterates for $T_{\text{max}}$ steps. At $t=0$, ${s_{q}}\left(0\right)$ is obtained using (16) and then is input into the policy network. In the early stages of training, actions are randomly sampled from $U(-1,1)$ to explore the environment without relying on the policy network output. Later, actions are determined by the mean and variance output of the policy network. The action distribution follows a multidimensional Gaussian distribution and is scaled within $\left({-1,1}\right)$. Then, based on the action distribution ${\pi_{\rho}}\left({\left.{{a_{q}}\left(0\right)}\right|{s_{q}}\left(0\right)}\right)$ with mean ${\mu_{\rho}}\left(0\right)$ and variance ${\sigma_{\rho}}\left(0\right)$, actions are sampled. The BS agent takes action ${a_{q}}\left(0\right)$ and receives a reward $r\left(0\right)$ following (20). Subsequently, the tuple $\left({{s_{q}}\left(0\right),{a_{q}}\left(0\right),r\left(0\right),{s_{q}}\left(1\right)}\right)$ is stored in the replay buffer to complete the iteration for $t=0$. This process can be iteratively repeated. The algorithm progresses to the next step at $t=1$. Here, the state transition moves from ${s_{q}}\left(0\right)$ to ${s_{q}}\left(1\right)$. When the step count reaches the maximum value $T_{\text{max}}$, the episode counter is incremented by $1$. If the number of samples in the experience replay buffer reaches a threshold, training is performed. The detailed update process is as follows.
Randomly select a batch of $bs$ tuples from the replay buffer. Input the current state ${s_{q}}\left(i\right)$ and the next state ${s_{q}}\left({i+1}\right)$ into the policy network, one can obtain the current action ${a_{q}}\left(i\right)$, log probabilities of actions $log\left({{\pi_{\rho}}\left({{a_{q}}\left(i\right)|{s_{q}}\left(i\right)}\right)}\right)$ with its mean and log standard deviation, the next time step actions ${a_{q}}\left({i+1}\right)$ and the log probabilities of the next time step actions $log\left({{\pi_{\rho}}\left({{a_{q}}\left({i+1}\right)|{s_{q}}\left({i+1}\right)}\right)}\right)$. The actions $a_{{}_{q}}^{bs}$ in a batch $bs$ are normalized using the following formula:
@@
The Q network parameters ${\theta_{1}}$ and ${\theta_{1}}$ are updatedA using gradient descent (GD) to minimize ${q_{los{s_{1}}}}$ and ${q_{los{s_{2}}}}$, which can be expressed as:
@@
$\tau$ is the soft update parameter, a fraction between 0 and 1 that controls the update speed of the target Q network parameters. So, the above four parameters and two target networks ${\bar{\theta}_{1}}$ and ${\bar{\theta}_{2}}$ are updated to their latest values. Then, the process continues to the next episode until $episode\_\max$ episodes are completed, obtaining the optimal parameters for the policy network ${\rho^{*}}$.
@@ III-B2 Testing stage
During the testing phase, the first step involves loading a pre-trained SSS model, which is then utilized to execute tests and derive the optimal policy ${\pi^{*}}\left({{a_{q}}\left(t\right)|{s_{q}}\left(t\right)}\right)$ using optimized parameters ${\rho^{*}}$. Similar to the training process, the initialization of the environment is required. During the testing phase, a loop is executed $episode_{t}est$ times to obtain the optimal channel allocation, transmission power, and the number of transmitted semantic symbols. Performance metrics, such as HSSE and SRS in V2V links, are computed and recorded. The pseudo-code for the testing phase is provided in Algorithm 2.
@@ IV-B Performance Evaluation
  • Semantic-aware Spectrum Sharing Algorithm based on deep deterministic policy gradient (DDPG) DRL: This baseline utilizes a DDPG DRL algorithm for spectrum sharing considering semantic data. In our comparisons, it is referred to as $ddpg\_with\_sc$.
  • Bits Spectrum Sharing Algorithm based on SAC DRL: In this baseline, actions involve selecting V2V links and determining the transmission power for the selected link, without considering the selection of semantic symbol lengths for each vehicle. We denote the baseline as $SAC\_without\_sc$.
  • Bits Spectrum Sharing Algorithm based on DDQN DRL: Similar to the second baseline, the difference lies in the utilization of DDQN DRL. We donate the baseline as $ddqn\_without\_sc$.
  • Bits Spectrum Sharing Algorithm based on Random selection: At each time step, BS acting as agents randomly selects shared V2V links and transmission power for the selected link. We donate the baseline as $rand\_without\_sc$.
@@
Fig.  4 illustrates the performance of three algorithms over time, measured in milliseconds (ms). The demand represents the successfully delivered payload by each algorithm. Here, semantic information has been converted to bits using a transforming factor ${u_{q}}$, resulting in a payload size of $25\times 1060$ bits. We observe that our proposed SSS algorithm and the $ddqn\_without\_sc$ successfully deliver the effective payload. In contrast, the $rand\_without\_sc$ fails to achieve this in the Fig. 4(c), due to its lack of a learning mechanism to dynamically adjust decisions based on the environment and tasks. Furthermore, the $ddqn\_without\_sc$ exhibits a noticeably slower descent rate compared to our proposed SSS algorithm. The successful payload delivery time varies for each vehicle in the $ddqn\_without\_sc$, as seen in the Fig.  4(a) and Fig.  4(b). Our SSS algorithm achieves effective delivery of semantic information in less than 5 milliseconds. This success can be attributed to delivering semantic information containing more valuable data compared to traditional equal-bit information. Secondly, the SSS method effectively utilizes semantic information, enabling more precise spectrum sharing. Consequently, each vehicle’s payload can rapidly descend and successfully complete delivery.
@@
Fig. 6 shows how average HSSE varies with increasing semantic demand size. The transform factor ${u_{q}}$ is set to 20, corresponding to $V2V$ semantic demand sizes ranging from $\left[{1,2,\cdots}\right]\times 530$ suts, with a transmission power of 23 $dBm$ for the V2I link transmitter. A consistent trend is observed: HSSE decreases with an increase in demand size, regardless of the algorithm used. Two main factors contribute to this trend. Firstly, Longer transmission times due to increased semantic demand sizes lead to prolonged interference of V2V links with V2I, resulting in reduced HSSE. Secondly, the increase in semantic demand sizes signifies a greater need for semantic information. Our SSS algorithm demonstrates notably higher HSSE and transmitting semantic information rather than bits. Semantic information conveys more meaningful and user-oriented data, enabling our SSS algorithm to more effectively handle the transmission of semantic demand sizes, thereby reducing interference from V2V links to V2I links and enhancing HSSE. Furthermore, transmitting semantic information contains richer information than transmitting traditional bits at the same HSSE. Notably, $rand\_without\_sc$ performs the worst among the three traditional spectrum sharing algorithms when the semantic payload size is less than $20\times 530$ suts. Surprisingly, in cases where the semantic payload significantly exceeds $20\times 530$, $rand\_without\_sc$ outperforms $ddqn\_without\_sc$. This is because in high-speed mobile IoV environments, $rand\_without\_sc$ exhibits a degree of flexibility, adapting better to dynamic environmental changes, while the $ddqn\_without\_sc$ may require more time to adapt to such variations.
@@
Fig. 7 shows the relationship between semantic demand size and SRS. We still set the transform factor ${u_{q}}$ as 20 and the transmission power of the V2I link transmitter as 23 $dBm$. As semantic demand sizes increase, SRS decreases accordingly. This reduction is due to the growing demand size, which increases the need for semantic information transmission within a unit of time, resulting in a decline in the SRS. Our SSS method outperforms the other four algorithms, exhibiting a slower decline. This advantage arises from our algorithm’s use of compressed semantic information rather than traditional bits. Due to the relatively smaller size of semantic information, it occupies fewer resources, enabling the maintenance of a higher V2V SRS. Beyond a semantic payload of $20\times 530$ suts, the V2V transmission probability significantly decreases for $ddqn\_without\_sc$. This is because $ddqn\_without\_sc$ struggles to effectively adapt to the dynamic and fast-paced environment of vehicular networks when dealing with a large amount of semantic payload. With increasing semantic payload, traditional bits transmission methods may lead to a sharp decline in transmission efficiency. With small semantic payloads, $rand\_without\_sc$ performs well, possibly due to its higher randomness. However, as the semantic demand size increases, our proposed algorithm demonstrates more stable and superior performance. This is attributed to the compression and transmission of semantic information, allowing our algorithm to more efficiently utilize limited spectrum resources and consequently improve the SRS.
@@
Fig. 8 shows the relationship between V2I link transmitter power and HSSE. We set the transform factor ${u_{q}}$ as 20 and the semantic demand size as $25\times 530{\rm{}}$ suts. Overall, as V2I power increases, average HSSE first rises and then stabilizes, indicating an upper limit to average HSSE with increasing V2I power. With higher transmission power, the V2I link can more effectively transmit semantic payloads, thereby enhancing HSSE. However, as power gradually increases, the transmission of semantic information becomes fixed, reaching a constant value after completing semantic information transmission. Our proposed algorithm consistently achieves higher HSSE across the entire power range compared to $ddqn\_without\_sc$, $rand\_without\_sc$, and $SAC\_without\_sc$, which transmit bits. Algorithms transmitting semantic information achieve higher HSSE. At low power, traditional algorithms exhibit a faster growth trend than algorithms transmitting semantic information. This is because traditional algorithms, focusing solely on the transmission of information rather than handling more complex semantic payloads, experience a rapid but still lower increase compared to our proposed algorithm.
@@
Fig. 10 shows the relationship between the number of vehicles and HSSE, indicating an overall increase with the growth of vehicle count. In the early stages, there is a rapid growth, followed by a slower increase. In the later stages, ultimately stabilizing. Our proposed algorithm outperforms $ddqn\_without\_sc$ and $rand\_without\_sc$. This is because our algorithm transmits semantic information, allowing flexible adaptation to diverse semantic requirements and enhancing HSSE. Additionally, the SAC reinforcement learning algorithm continuously improves decision quality by accumulating experience, adapting to complex communication environments. Initially, as the number of vehicles increases, the system experiences improved communication efficiency. However, continued growth saturates system resources, leading to competition and conflicts among communication demands. This phenomenon could be attributed to potential bottlenecks in communication resources and spectrum availability, limiting the system’s ability to maintain the same rate of performance improvement in large-scale scenarios.

### v4

@@ II-C Optimization Problem
${P_{0}}:\mathop{\max}\limits_{{b_{q}},{P_{q}},{u_{q}}}\sum\limits_{w=1}^{W}{HSSE_{q,w}^{VI}}+\sum\limits_{w=1}^{W}{\sum\limits_{q=1}^{Q}{{b_{q}}\left[w\right]}}p_{q}^{VV}\left[w\right]$ (15a)
@@
Constraints (15b), (15c) and (15d) represent the channel allocation. (15c) states that each V2V link can only share one of the V2I links, (15d) specifies that each V2I link can only be allocated to one V2V link, (15e) constrains the average number of semantic symbols for each word within a certain range, and (15f) constrains the minimum value of semantic similarity. The physical meaning of this constraint is to ensure that the semantic content of the transmitted message maintains a predefined level of accuracy, guaranteeing that the essential information conveyed by the message is both comprehensible and relevant.
@@ III-A2 Action
The key actions include channel selection for V2V links, transmission power control, and the selection of the length of semantic symbols by the V2V transmitter vehicle in the shared V2I link. The agent’s actions can be represented as a vector containing multiple elements. Specifically, it can be represented as shown in (17):
${a_{q}}\left(t\right)=\left({{b_{q}\left(t\right)},{P_{q}}\left(t\right),{u_{q}}\left(t\right)}\right),$ (17)
where ${b_{q}}\left(t\right)$ represents the channel selection, ${P_{q}}\left(t\right)$ represents the transmission power chosen by the V2V link transmitter vehicles, and ${u_{q}}\left(t\right)$ represents the average number of semantic symbols used for each word which means the length of semantic symbols chosen by the V2V transmitter vehicle in the shared V2I link. These actions’ values are usually confined within the range $\left[{-1,0}\right]$, which has been selected for efficient learning and convergence. Subsequently, these values undergo a linear mapping process to obtain actual parameters or lengths. Defining this action space enables flexible decision-making in semantically-aware spectrum sharing within high-speed mobile IoV environments, aiming to maximize the HSSE and SRS over V2V links.
@@ III-B SSS Algorithm
The SSS algorithm regarding the training phase and is illustrated in Algorithm 1 and is described in detail as follows, including how it is deployed in practical IoV networks.
@@ III-B1 Training stage
where $sh{a_{new}}$ and $sh{a_{old}}$ is the shadow fading at the new and old position, respectively. $X\sim N(0,{3^{2}})$ , ${\Delta d}$ represents the distance between paths. ${{d_{dec}}}$ is the decorrelation distance used to describe the correlation of shadow fading. All the aforementioned models and parameters comply with the C-V2X 3GPP standard communication protocol, ensuring interoperability across various IoV platforms and devices [48].
@@
Next, initialize the size of the semantic data packets to be transmitted, the specified transmission time, and the link. The algorithm iterates for $T_{\text{max}}$ steps. For each time step, the cloud server receives the state observations from the vehicles. At $t=0$, ${s_{q}}\left(0\right)$ is obtained using (16) and then is input into the policy network. In the early stages of training, actions are randomly sampled from $U(-1,1)$ to explore the environment without relying on the policy network output. Later, actions are determined by the mean and variance output of the policy network. The action distribution follows a multidimensional Gaussian distribution and is scaled within $\left({-1,1}\right)$. Then, based on the action distribution ${\pi_{\rho}}\left({\left.{{a_{q}}\left(0\right)}\right|{s_{q}}\left(0\right)}\right)$ with mean ${\mu_{\rho}}\left(0\right)$ and variance ${\sigma_{\rho}}\left(0\right)$, actions are sampled. The BS agent takes action ${a_{q}}\left(0\right)$ and receives a reward $r\left(0\right)$ following (20). Subsequently, the tuple $\left({{s_{q}}\left(0\right),{a_{q}}\left(0\right),r\left(0\right),{s_{q}}\left(1\right)}\right)$ is stored in the replay buffer by the cloud sever to complete the iteration for $t=0$. This process can be iteratively repeated. The algorithm progresses to the next step at $t=1$. Here, the state transition moves from ${s_{q}}\left(0\right)$ to ${s_{q}}\left(1\right)$. If the number of samples in the experience replay buffer reaches a threshold,the server randomly selects a batch of $bs$ tuples from the replay buffer. Input the current state ${s_{q}}\left(i\right)$ and the next state ${s_{q}}\left({i+1}\right)$ into the policy network, one can obtain the current action ${a_{q}}\left(i\right)$, log probabilities of actions $log\left({{\pi_{\rho}}\left({{a_{q}}\left(i\right)|{s_{q}}\left(i\right)}\right)}\right)$ with its mean and log standard deviation, the next time step actions ${a_{q}}\left({i+1}\right)$ and the log probabilities of the next time step actions $log\left({{\pi_{\rho}}\left({{a_{q}}\left({i+1}\right)|{s_{q}}\left({i+1}\right)}\right)}\right)$. The actions $a_{{}_{q}}^{bs}$ in a batch $bs$ are normalized by the cloud server:
@@
The Q network parameters ${\theta_{1}}$ and ${\theta_{1}}$ are updated using gradient descent (GD) to minimize ${q_{los{s_{1}}}}$ and ${q_{los{s_{2}}}}$, which can be expressed as:
@@
$\tau$ is the soft update parameter, a fraction between 0 and 1 that controls the update speed of the target Q network parameters. So, the above four parameters and two target networks ${\bar{\theta}_{1}}$ and ${\bar{\theta}_{2}}$ are updated to their latest values. Then, the process continues to the next episode until $episode\_\max$ episodes are completed, obtaining the optimal parameters for the policy network ${\rho^{*}}$. The Q networks and the policy network and their related parameters are all updated by the cloud sever. After training, the cloud server broadcasts the updated model parameters to the BS, which then disseminates them to all vehicles in the network.
@@ III-B2 Testing stage
During the testing phase, the first step involves loading a pre-trained SSS model, which is then utilized to execute tests and derive the optimal policy ${\pi^{*}}\left({{a_{q}}\left(t\right)|{s_{q}}\left(t\right)}\right)$ using optimized parameters ${\rho^{*}}$. Similar to the training process, the initialization of the environment is required by the cloud server. During the testing phase, a loop is executed $episode_{t}est$ times to obtain the optimal channel allocation, transmission power, and the number of transmitted semantic symbols. Performance metrics, such as HSSE and SRS in V2V links, are computed and recorded. The pseudo-code for the testing phase is provided in Algorithm 2.
@@ IV-B Performance Evaluation
  • Semantic-aware Spectrum Sharing Algorithm based on DDPG DRL: This baseline utilizes a DDPG DRL algorithm for spectrum sharing considering semantic data. In our comparisons, it is referred to as $DDPG\_with\_sc$.
  • Bits Spectrum Sharing Algorithm based on SAC DRL [50]: In this baseline, actions involve selecting V2V links and determining the transmission power for the selected link, without considering the selection of semantic symbol lengths for each vehicle. We denote the baseline as $SAC\_without\_sc$.
  • Bits Spectrum Sharing Algorithm based on DDQN DRL [26]: Similar to the second baseline, the difference lies in the utilization of DDQN DRL. We donate the baseline as $DDQN\_without\_sc$
  • Bits Spectrum Sharing Algorithm based on Random selection [27]: At each time step, BS acting as agents randomly selects shared V2V links and transmission power for the selected link. We donate the baseline as $rand\_without\_sc$.
@@
Fig. 4 illustrates the performance of three algorithms over time, measured in 100 milliseconds (ms). The demand represents the successfully delivered payload by each algorithm. Semantic information is converted to bits using a transforming factor ${u_{q}}$, resulting in a payload size of $25\times 1060$ bits. We observe the success rates of the SSS algorithm and $DDQN\_without\_sc$ are 100%, compared to 75% with $rand\_without\_sc$ due to its lack of a learning mechanism to dynamically adjust decisions. The average response time of our SSS algorithm is 3.8 milliseconds, while $DDQN\_without\_sc$ algorithm has an average response time of 14 milliseconds, representing a 72.9% reduction in latency. This improvement is primarily due to the SSS algorithm’s ability to leverage semantic information to deliver more valuable data and reduce redundancy, thus lowing transmission delay. Additionally, the algorithm’s real-time decision-making capabilities and efficient resource allocation quickly adapt to changing conditions, ensuring low latency and high throughput. Low latency is also indicative of enhanced user experience, as it ensures timely and efficient communication, which is crucial for applications in IoV environments [51]. The successful payload delivery time varies for each vehicle in the $DDQN\_without\_sc$, as seen in the Fig. 4(a) and Fig. 4(b).
@@
Fig. 6 shows how average HSSE varies with increasing semantic demand size. The transform factor ${u_{q}}$ is set to 20, corresponding to $V2V$ semantic demand sizes ranging from $\left[{1,2,\cdots}\right]\times 530$ suts, with a transmission power of 23 $dBm$ for the V2I link transmitter. HSSE decreases with an increasing demand size, regardless of the algorithm used, which is because that increased semantic demand sizes lead to prolonged interference of V2V links with V2I, resulting in reduced HSSE and increased semantic demand sizes signifies a greater need for semantic information. Our SSS algorithm demonstrates notably higher HSSE which means the lower energy consumption and transmitting semantic information rather than bits. Semantic information conveys more meaningful and user-oriented data, enabling our SSS algorithm to more effectively handle the transmission of semantic demand sizes, thereby reducing interference from V2V links to V2I links and enhancing HSSE. Furthermore, transmitting semantic information contains richer information than transmitting traditional bits at the same HSSE. Notably, $rand\_without\_sc$ performs better than outperforms $DDQN_{w}ithout_{s}c$ when the semantic payload size is more than $20\times 530$ suts. This is because in high-speed mobile IoV environments, $rand\_without\_sc$ exhibits a degree of flexibility, adapting better to dynamic environmental changes, while the $DDQN\_without\_sc$ may require more time to adapt to such variations. Furthermore, HSSE can also reflect user experience, as it indicates the efficiency and effectiveness of semantic information transmission.
@@
Fig. 7 shows the relationship between semantic demand size and SRS. ${u_{q}}$ is set as 20 and transmission power of the V2I link transmitter as 23 $dBm$. As semantic demand sizes increase, SRS decreases. This reduction occurs because a higher demand size requires more semantic information transmission within a given time, leading to a decline in SRS. Our SSS method exhibits a slower decline compared to the other algorithms which means the lower speed of energy consumption. We attribute to the use of compressed semantic information instead of traditional bits. The smaller size of semantic information reduces resource usage, maintaining higher V2V SRS. Beyond a semantic payload of $20\times 530$ suts, the V2V transmission probability significantly decreases for $DDQN\_without\_sc$. This is because this method struggles to effectively adapt to the dynamic vehicular networks when dealing with a large amount of semantic payload. With increasing semantic payload, traditional bits transmission methods may lead to a sharp decline in transmission efficiency. In contrast, $rand\_without\_sc$ performs relatively well with smaller payloads due to its randomness but declines with larger payloads. Our proposed algorithm, by compressing and efficiently transmitting semantic information, demonstrates more stable and superior performance, effectively utilizing limited spectrum resources and improving SRS.
@@
Fig. 8 shows the relationship between V2I link transmitter power and HSSE. ${u_{q}}$ is set as 20 and semantic demand size as $25\times 530{\rm{}}$ suts. As V2I power increases, initially rises and then stabilizes, indicating an upper limit to HSSE with increasing power. Higher transmission power enables V2I link TO more effectively transmit semantic payloads, thereby enhancing HSSE. However, beyond a certain point, further increases in power do not significantly impact HSSE, as the transmission of semantic information reaches a constant value. Our proposed algorithm consistently achieves higher HSSE which means the lower energy consumption across the entire power range compared to $DDQN\_without\_sc$, $rand\_without\_sc$, and $SAC\_without\_sc$, which transmit bits. Algorithms transmitting semantic information achieve higher HSSE. At low power levels, traditional algorithms exhibit a faster growth trend than algorithms transmitting semantic information. This is because traditional algorithms, focusing solely on the transmission of information rather than handling more complex semantic payloads, experience a rapid but still lower increase compared to our proposed algorithm.
@@
Fig. 10 shows the relationship between the number of vehicles and HSSE, showing an overall increase as the vehicle count grows. Our proposed algorithm outperforms other $without\_sc$ methods. As the number of vehicles increases, more users compete for channel resources, causing HSSE to initially rise. However, as vehicles’ number still growing, this also intensifies competition for limited spectrum resources, leading to potential bottlenecks in communication resources and spectrum availability. Our SSS excels in addressing this bottleneck by employing semantic information-based dynamic spectrum management. The DRL-based methods $without\_sc$ still outperform traditional random algorithms due to the optimized decision-making processes.